9. 100 mL of 0.5M HNO2 (Ka = 4.5*10-4) is titrated with 0.5M KOH. What is the pH after 100mL of 0.5M KOH is titrated into the acid? (12)
Answer #1

Givens Hno, tkoH KNO, tho ncmoles) molants a la X1000 looml of a.5 mm Hoog 0-5 X 100 no. of moles of Hoop = 0 100= 0.05 moles2.22X10 0.25 a²=5-5501012 2 = 2.36.816 2.36X10 [OH-] = 2-368106 -log(04-) = -log (2.36x16] pOH = 5.6% ph + polt = 14 .. ph14